# Computation of Magnetohydrodynamic Equilibria with Voigt Regularization

### v3

@@ I Introduction
The approach of Voigt-MHD bears some similarities with existing MHS solvers that do not assume nested flux surfaces. For example, the HINT2 code alternately relaxes the pressure and the magnetic field in a way that mimics time evolution.[22, 23] In the pressure relaxation step, the new pressure is obtained at each point by averaging the current pressure along the field line over a given length. The new pressure is then used in the magnetic field relaxation step, which solves the momentum and induction equations of the resistive MHD system, holding the pressure fixed. These steps continue until an equilibrium is reached. In an alternative approach, the SIESTA code directly steps the adiabatic pressure equation and the induction equation simultaneously.[24] However, a small resistivity has to be periodically applied to allow breaking of magnetic surfaces. The Voigt-MHD approach differs from these methods by being grounded in a well-defined system of partial differential equations rather than a prescription of somewaht ad hoc procedures.
@@ II.1 Governing Equations and Boundary Conditions
The non-dimensional form of Voigt-regularized, incompressible MHD system is given by
@@ III.2 Ideal MHD Equilibria
Then we perturb the field by multiplying $\psi_{0}$ by a factor of $1+\epsilon\cos\left(2\pi x/L_{x}\right)$, where $\epsilon$ is a small parameter. This perturbation seeds a narrow island in the field.
@@
Inspecting the residual force distribution reveals that the unbalanced force is mostly localized along the separatrix near the X-point, where the current density $J$ becomes spiky (Figure 2(a)). Inspecting the current density profile also suggests that this numerical calculation does not have sufficient resolution to resolve the spiky current density. Moreover, because the force balance condition $-\nabla p+\boldsymbol{J}\times\boldsymbol{B}=0$ in 2D implies $\boldsymbol{B}\cdot\nabla J=0$, the spiky $J$ near the X-point should spread out exactly along the separatrix such that $J=J\left(\psi\right)$. This condition is clearly not satisfied in Figure 2(a). Figure 2(b) further reveals the difficulty, showing that the residual force is mostly concentrated along the separatrix. To resolve this thin current structure, we must have high resolution not only near the X-point, but also along the separatrix. However, even if we have sufficient numerical resolution, the spreading of current density along the separatrix would likely take a long time.
@@ III.3 Resistive MHD Equilibria with External Electric Field
corresponding to a uniformly sheared magnetic field $\boldsymbol{B}=y\boldsymbol{\hat{x}}$ with an added perturbation. We set $L_{x}=7$, $L_{y}=1,$ and $\epsilon=0.1$. The resolution is $N_{x}=N_{y}=128$. The viscosity and resistivity are set to $\eta=\nu=10^{-4}.$ We vary the Voigt coefficients for the following values $\alpha_{1}=\alpha_{2}=0,$ $0.1$, $0.2$, and $0.5$. The corresponding time-steps are $\Delta t=0.05$, $2$, $5$, and $30$. In this example, the plasma velocity asymptotically goes to zero. Therefore, the CFL condition is only limited by waves, and we can use much larger time-steps with the Voigt regularization.

### v4

@@ I Introduction
The approach of Voigt-MHD bears some similarities with existing MHS solvers that do not assume nested flux surfaces. For example, the HINT2 code alternately relaxes the pressure and the magnetic field in a way that mimics time evolution.[22, 23] In the pressure relaxation step, the new pressure is obtained at each point by averaging the current pressure along the field line over a given length. The new pressure is then used in the magnetic field relaxation step, which solves the momentum and induction equations of the resistive MHD system, holding the pressure fixed. These steps continue until an equilibrium is reached. In an alternative approach, the SIESTA code directly steps the adiabatic pressure equation and the induction equation simultaneously.[24] However, a small resistivity has to be periodically applied to allow breaking of magnetic surfaces. The Voigt-MHD approach differs from these methods by being grounded in a well-defined system of partial differential equations rather than a prescription of somewhat ad hoc procedures.
@@ II.1 Governing Equations and Boundary Conditions
The non-dimensional form of the Voigt-regularized, incompressible MHD system is given by
@@ III.2 Ideal MHD Equilibria
We perturb the field by first adding $-\epsilon\left(L_{x}/2\pi\right)\cos\left(2\pi x/L_{x}\right)\cos\left(\pi y/L_{y}\right)$ to $\psi_{0}$, then multiplying the resulting flux function by a factor of $1+\epsilon\cos\left(2\pi x/L_{x}\right)$, where $\epsilon$ is a small parameter. This perturbation seeds a narrow island in the field.
@@
The residual force distribution [Fig. 2(b)] reveals that the unbalanced force is mostly localized along the separatrix near the X-point, where the current density $J$ becomes spiky [Fig. 2(a)]. Inspecting the current density profile also suggests that this numerical calculation does not have sufficient resolution to resolve the spiky current density. Moreover, because the force balance condition $-\nabla p+\boldsymbol{J}\times\boldsymbol{B}=0$ in 2D implies $\boldsymbol{B}\cdot\nabla J=0$, the spiky $J$ near the X-point should spread out exactly along the separatrix such that $J=J\left(\psi\right)$. This condition is clearly not satisfied in Figure 2(a). To resolve this thin current structure, we must have high resolution not only near the X-point, but also along the separatrix. However, even if we have sufficient numerical resolution, the spreading of current density along the separatrix would likely take a long time.
@@ III.3 Resistive MHD Equilibria with External Electric Field
corresponding to a uniformly sheared magnetic field $\boldsymbol{B}=y\boldsymbol{\hat{x}}$ with an added perturbation. We set $L_{x}=7$, $L_{y}=1,$ and $\epsilon=0.1$. The resolution is $N_{x}=N_{y}=128$. The viscosity and resistivity are set to $\eta=\nu=10^{-4}.$ We vary the Voigt coefficients for the following values $\alpha_{1}=\alpha_{2}=0,$ $0.1$, $0.2$, and $0.5$. The corresponding time-steps are $\Delta t=0.05$, $2$, $5$, and $30$. In this example, the plasma velocity asymptotically goes to zero. Therefore, the stability condition is only limited by waves, and we can use much larger time-steps with the Voigt regularization.